# QRMeM: Unleash the Length Limitation through Question then Reflection Memory Mechanism

### v1

@@ Abstract
While large language models (LLMs) have made notable advancements in natural language processing, they continue to struggle with processing extensive text. Memory mechanism offers a flexible solution for managing long contexts, utilizing techniques such as compression, summarization, and structuring to facilitate nuanced and efficient handling of large volumes of text. However, existing techniques face challenges with static knowledge integration, leading to insufficient adaptation to task-specific needs and missing multi-segmentation relationships, which hinders the dynamic reorganization and logical combination of relevant segments during the response process. To address these issues, we introduce a novel strategy, Question then Reflection Memory Mechanism (QRMeM), incorporating a dual-structured memory pool. This pool synergizes static textual content with structured graph guidance, fostering a reflective trial-and-error approach for navigating and identifying relevant segments. Our evaluation across multiple-choice questions (MCQ) and multi-document question answering (Multi-doc QA) benchmarks showcases QRMeM’s enhanced performance compared to existing approaches. [^1] The code used in our work will be released soon..
@@
Bo Wang1 , Heyan Huang1, Yixin Cao 2 , Jiahao Ying2, Wei Tang 4, Chong Feng 1 1Beijing Institute of Technology, China 2Fudan University, China 3Singapore Management University, Singapore 4 University of Science and Technology of China, China bwang@bit.edu.cn
@@ 1 Introduction
Figure 1: Question then Reflection Memory mechanism. This memory is composed of structured memory and long-term memory, which collaborate during navigation. During iterative navigation, the structured graph guides the selection of relevant segments. QRMeM allows for the exploration of segment combinations based on experience throughout the iterative navigation.
Retrieval-based methods aim to filter relevant parts from the original long documents Zhang et al. (2023); Yu et al. (2023c). They achieve this through various retrieval operations, such as dense, sparse, and reranking, between the text and the query. Another line of approaches, known as memory mechanism methods Liang et al. (2023), takes the source as a memory pool. These methods offer a solution for managing long contexts and provide an opportunity to handle extraneous documents effectively. Existing memory mechanisms employ several strategies for indexing extensive information. The strategies include storing most recentest acquired information Liang et al. (2023); Ebbinghaus (2013), selecting contents based on their relevance, importance, and topics Zhong et al. (2024); Park et al. (2023), and handling long contexts through techniques such as compression Pan et al. (2024), summarization Chen et al. (2023a), and structuring Hu et al. (2023a).
@@ 2.1 Dual-structure Memory Construction
The approach involves partitioning the memory component into two distinct sections, structured and long-term static memory. The construction is depicted in Figure 2. The structured aspects of the memory resemble human memory anchors, assisting in finding more relevant chunks related to the query as one’s understanding deepens. The long-term portion serves as a reference preventing information loss or misinterpretation.
@@
Doubt serves as a fundamental catalyst for the acquisition of knowledge. We advance the methodology of knowledge graph updates through a mechanism predicated on question generation. This mechanism is formalized as ${\mathbb{Q}}_{k}=\text{LLM}(T,{\mathbf{S}}_{k},{\mathbf{E}}_{k},{\mathbf{R}}_{k})$, where ${\mathbb{Q}}_{k}$ denotes the generated question leveraging a Large Language Model (LLM) and the $T$ represents the summary of the entire document. To ensure the diversity of the generated questions, a novel question is incorporated into the question pool only if its ROUGE-L similarity score with any existing question is less than 0.6.333Excessive number of questions could introduce further noise, we empirically selected 0.6 as threshold. Entities and relations are further supplemented through similar techniques based on the questions ${\mathbb{Q}}_{k}$ during the initial stage.
Global Graph Combination: In the process of integrating sub-graphs $\{G_{1},G_{2},...,G_{{\textnormal{K}}}\}$ into a comprehensive global graph $G$, we further apply the entity disambiguation and relation merging. After entity disambiguation, entities associated with distinct segments are further elaborated into triples of the form ${e}_{k,j}=\{{e}^{{}^{\prime}}_{k,j},{\mathbb{M}}_{k,j},{\mathbb{N}}\}$, thereby facilitating the update of paragraph indices. Here, ${\mathbb{M}}_{k,j}$ and ${\mathbb{N}}$ represent the sets of entity mentions and segment indices, respectively. after merging entity pairs ($e_{n,i},e_{k,i^{\prime}}$) and ($e_{n,j},e_{k,j^{\prime}}$) into consolidated entities $e_{(n,k),i}$ and $e_{(n,k),j}$, Following the initial setup phase, we generate questions for merging relationships $Q_{s}$ according to the entities and their original segments, which could benefit from rethinking the relation between entity pairs. Consequently, relations $r_{n,(i,j)}$ and $r_{k,(i^{\prime},j^{\prime})}$ are amalgamated into a unified relation $r_{(k,n),(i,j)}=\text{LLM}(r_{n,(i,j)},{\mathbf{S}}_{n},r_{k,(i^{\prime},j^{\prime})},{\mathbf{S}}_{k},Q_{s},Q,T)$. Static segments will be associated with the entity nodes, thereby organizing the connection between the structured and long-term parts in the memory pool, as shown in Figure 2.
2.2 Reflection-based Relevant Segments Navigation
@@
Graph Expansion Search. As the number of nodes increases, the model’s ability to filter nodes and the window length both fall short of meeting the need of Entity Trial. Graph Expansion Search further leverages the structured part of the memory pool, $G$, to enhance exploration from segments. $\mathbb{E}_{s}$ is first expanded by incorporating adjacent nodes within the graph. The expansion is guided by the computation of similarity scores between the given query $Q$ and the $\mathbb{R}_{s}$, retaining nodes $\mathbb{E}^{\prime}$ with high similarity. The entity set $\mathbb{E}$ is updated through $\mathbb{E}\leftarrow\mathbb{E}^{\prime}\cup\mathbb{E}$ following each expansion iteration. The expansion process continues until either a predefined threshold of expansion iterations is reached or no new edges with sufficiently high similarity scores are found. The resultant expanded set of entities $\mathbb{E}$ is then to generate the elaborated query through ${\mathbb{Q}}=LLM(\mathbb{E},\mathbb{R},Q)$. ${\mathbb{Q}}$ is utilized to retrieve relevant segments, which are constructed into $c$ to subsequently address $Q$.
@@
QRMeM. To trace the experience across multiple iterations, QRMeM utilizes a reflection-based navigation strategy that dynamically traverses the graph $G$, which is shown in Figure 3 Initially, QRMeM sets up an additional segmentation set, ${\mathbb{S}}_{\text{add}}$, as $\emptyset$ for further navigation. For each iteration, the real relevant segments ${\mathbb{S}}_{\text{mix}}$ is formed by combination of ${\mathbb{S}}_{\text{imp}}$ and ${\mathbb{S}}_{\text{add}}$. Then the LLM is employed to evaluate whether ${\mathbb{S}}_{\text{mix}}$ can answer the question $Q$ by the feedback of the question $Q$ and the reason $R$ for the failure is identified if not. While the ${\mathbb{S}}_{\text{mix}}$ is insufficient for the current query, the first-order adjacent nodes $\mathbb{E}_{\text{adj}}$ of the current entities and all connecting edges $\mathbb{R}$ are collected. Drawing upon the lessons learned from the previous iteration, the query, reason, and current entities $[Q;R;\mathbb{E}_{\text{s}}]$ are used to select the next expanded node. ContrieverIzacard et al. (2022) is used for calculating the similarity score between the $[Q;R;\mathbb{E}_{\text{s}}]$ and each edge in $\mathbb{R}$. After the most relevant adjust node $e$ is selected and merged with the new entity $e$ into the entity set $\mathbb{E}$, the segments corresponding to the new entity $e$ are collected. All segments not present in ${\mathbb{S}}_{\text{mix}}$ are added to ${\mathbb{S}}_{\text{add}}$ with the similarity score. Allowing for thorough updates to ${\mathbb{S}}_{\text{add}}$, a crucial step is to ensure that the context window remains sufficient after updates. ${\mathbb{S}}_{\text{add}}$ is processed and filtered according to the similarity score while adhering to the window constraints. This iterative segment combination, window checking, question evaluation, and sub-graph expansion process continues until the question $Q$ is satisfactorily answered. The detail of the algorithm is shown in Algorithm 1.
@@ 3.1 Baselines
To evaluate the effectiveness of the proposed method, we compare the QRMeM with several methods. Retrieval Methods. We use dense retrievers Contriever Izacard et al. (2022) and BM25 for all datasets. Vanilla, we use the model with 4096 context windows, including Llama-70b Touvron et al. (2023) and gpt-3.5-turbo. Summary, we summarize the long context to fit the context window size. DECOMFu et al. (2021) simplifies complex questions into sub-questions, answers them sequentially. LLMLingua Pan et al. (2024), compresses prompts to speed up inference and enhance response efficiency in large language models. Long context model, including Vicuna-7b-32k, Vicuna-13b-16k, Longchat-7b-32k, Longchat-13b-16k,Llama-70b-32k. MemWalker Chen et al. (2023a), which utilize a tree of summary nodes for long context understanding.
@@ 3.2 Experiment Setting
The official LLaMa-2-Chat and gpt-3.5-turbo are used for the majority of our experiments. The two models both have a maximum 4,096 token context length. We utilize the model in a zero-shot prompting fashion without further fine-tuning or using few-shot examples for our tasks. For both memory tree construction and generating action and reasoning for navigation, we use top-p sampling. The temperature is initially set to a default value, but if the output cannot be parsed, we increase the temperature and retry up to four times. The segment size is set to 600 for all datasets.
Qualtiy Hotpot 2Wiki Musique
@@ 3.4 Main Result
1) Our study presents significant advancements over retrieval methods on the Quality dataset, on both the LLaMA 70B and GPT-3.5-turbo models. Detailed results related to retrieval methods are referred to in Appendix B. Notably, our approach, when compared to models utilizing longer context windows, achieves results comparable to those of GPT-3.5-16K, thereby outperforming other models designed for long-text comprehension. This highlights the effectiveness of our method in selecting relevant segments, promoting a comprehensive understanding of complex documents.
2) Furthermore, our methods demonstrated exceptional performance on Multi-Doc QA datasets. Specifically, on the 2Wiki and Musique datasets, our approach achieved remarkable results and yields comparable results to existing retrieval models on the Hotpot dataset. Even though BM25 obtained a satisfactory result on Hotpot, it still struggles with complex tasks like Quality and Musique. Across all three datasets, our method outperformed long-document comprehension models. This indicates our approach’s adaptability to multi-hop tasks that require the understanding of entity relationships, capturing these relationships more effectively than long-document models. The lesser improvement on fewer inference steps Hotpot questions and more significant advancements on the complex Musique and Quality tasks illustrate our method’s effectiveness in organizing original paragraphs for complex tasks. The reason GPT-3.5-turbo outperforms GPT-3.5-turbo might be due to differences in their document understanding capabilities. GPT-3.5-turbo tends to focus more on the overall context, while GPT-3.5-turbo pays closer attention to details. This is also why GPT-3.5-16k shows better performance on QuALITY datasets.
3) Three strategies proposed in our study outperformed retrieval-enhanced and direct truncation strategies. Entity Trial’s performance improvement demonstrates the effectiveness of iterative exploration. By using entities as anchors and leveraging the reflective iterative exploration of segments, it shows the impact of reflective iteration in acquiring relevant segments. Graph Expansion Search, although lacking an iterative navigation process, still utilizes structured memory to guide the LLM. This indicates that our query-based structured memory provides better guidance and relevance compared to predefined indexing. Among all the methods, QRMeM achieves the best performance. It combines the strengths of both Entity Trial and Graph Expansion Search, using the graph to guide the model and leveraging past errors to inform future steps. This approach allows the model to utilize specific structured knowledge while retaining the ability to iteratively combine different segments.
@@ 3.7 The impact of irrelevant information
To further investigate the influence of irrelevant documents, we conducted additional tests on cross-document multi-hop questions. In this section, we retained the original wiki support text from the Long Bench dataset without chunking the long text, enabling us to more clearly identify the impact of irrelevant documents. Following Liu et al. (2023), we investigated the impact of varying the placement of supporting segments within a document, and the result is shown as 3 Specifically, we positioned the corresponding support segments at the beginning, middle, and tail of the overall document. Our findings are consistent with those reported by Liu et al. (2023), indicating that the placement of information can significantly affect the model’s performance on the Multi-Doc QA task. Therefore, in our subsequent experiments, we placed the multi-hop support text at both the beginning and end of the document, ensuring that the supporting segments remain in these positions to avoid any positional impact.
(a) Hotpot
@@
Irrelevant documents utilized to increase the distance are retained, represented as [$S_{1};S_{non};S_{2}$], where $S_{1}$ and $S_{2}$ are the supporting segments, and $S_{non}$ represent sequences of non-relevant segments. The outcomes are depicted in Figure 5. The results indicate that although the model exhibits heightened attention to the positions at the beginning and end of the text, its performance is still impacted by the distance between documents. Effectiveness consistently declines as distance increases on both the Hotpot and Musique datasets.

### v2

@@ Abstract
While large language models (LLMs) have made notable advancements in natural language processing, they continue to struggle with processing extensive text. Memory mechanism offers a flexible solution for managing long contexts, utilizing techniques such as compression, summarization, and structuring to facilitate nuanced and efficient handling of large volumes of text. However, existing techniques face challenges with static knowledge integration, leading to insufficient adaptation to task-specific needs and missing multi-segmentation relationships, which hinders the dynamic reorganization and logical combination of relevant segments during the response process. To address these issues, we introduce a novel strategy, Question then Reflection Memory Mechanism (QRMeM), incorporating a dual-structured memory pool. This pool synergizes static textual content with structured graph guidance, fostering a reflective trial-and-error approach for navigating and identifying relevant segments. Our evaluation across multiple-choice questions (MCQ) and multi-document question answering (Multi-doc QA) benchmarks showcases QRMeM’s enhanced performance compared to existing approaches. Our code is available at https://github.com/wang202111/qrmem.
@@
Bo Wang1,3 , Heyan Huang1,3, Yixin Cao2, Jiahao Ying4, Wei Tang5, Chong Feng1,3 1Beijing Institute of Technology, China 2Fudan University, China 3Southeast Academy of Information Technology, Beijing Institute of Technology, China 4Singapore Management University, Singapore 5 University of Science and Technology of China, China bwang@bit.edu.cn
@@ 1 Introduction
Retrieval-based methods aim to filter relevant parts from the original long documents Zhang et al. (2023); Yu et al. (2023c). They achieve this through various retrieval operations, such as dense, sparse, and reranking, between the text and the query.
Figure 1: Question then Reflection Memory mechanism is divided into two parts. The upper part, Structured Memory, is a graph constructed from entities and relations extracted from the text. The lower part, Original Segmentation, is the original segmentation. These two parts are linked by the co-occurrence relationships of entities appearing in the segments. QRMeM starts from the core entity of the question and expands on the Structured Memory to find the most relevant next entity node. It uses this to expand and revise the segment sets until the corresponding segment set can answer the question.
Another line of approaches, known as memory mechanism methods Liang et al. (2023), takes the source as a memory pool. These methods offer a solution for managing long contexts and provide an opportunity to handle extraneous documents effectively. Existing memory mechanisms employ several strategies for indexing extensive information. The strategies include storing most recentest acquired information Liang et al. (2023); Ebbinghaus (2013), selecting contents based on their relevance, importance, and topics Zhong et al. (2024); Park et al. (2023), and handling long contexts through techniques such as compression Pan et al. (2024), summarization Chen et al. (2023a), and structuring Hu et al. (2023a).
@@ 2.1 Dual-structure Memory Construction
The approach involves partitioning the memory component into two distinct sections, structured and long-term static memory. The construction is depicted in Figure 2. The structured aspects of the memory resemble human memory anchors, assisting in finding more relevant chunks related to the query as one’s understanding deepens. The Structured Memory serves only as guidance to avoid error accumulation during the graph construction process, while the Original Segmentation is used to answer questions, further preventing information loss or misinterpretation.
@@
Doubt serves as a fundamental catalyst for the acquisition of knowledge. We advance the methodology of knowledge graph updates through a mechanism predicated on question generation. This mechanism is formalized as ${\mathbb{Q}}_{k}=\text{LLM}(T,{\mathbf{S}}_{k},{\mathbf{E}}_{k},{\mathbf{R}}_{k})$, where ${\mathbb{Q}}_{k}$ denotes the generated question leveraging a Large Language Model (LLM) and the $T$ represents the summary of the entire document. To ensure the diversity of the generated questions, a novel question is incorporated into the question pool only if its ROUGE-L similarity score with any existing question is less than 0.6.222Excessive number of questions could introduce further noise, we empirically selected 0.6 as threshold. Entities and relations are further supplemented through similar techniques based on the questions ${\mathbb{Q}}_{k}$ during the initial stage.
Global Graph Combination: In the process of integrating sub-graphs $\{G_{1},G_{2},...,G_{{\textnormal{K}}}\}$ into a comprehensive global graph $G$, we further apply the entity disambiguation and relation merging. After entity disambiguation, entities associated with distinct segments are further elaborated into triples of the form ${e}_{k,j}=\{{e}^{{}^{\prime}}_{k,j},{\mathbb{M}}_{k,j},{\mathbb{N}}_{k,j}\}$, thereby facilitating the update of paragraph indices. Here, ${\mathbb{M}}_{k,j}$ and ${\mathbb{N}}$ represent the sets of entity mentions and segment indices, respectively. After merging entity pairs ($e_{n,i},e_{k,i^{\prime}}$) and ($e_{n,j},e_{k,j^{\prime}}$) into consolidated entities $e_{(n,k),i}$ and $e_{(n,k),j}$, Following the initial setup phase, we generate questions for merging relationships $Q_{s}$ according to the entities and their original segments, which could benefit from rethinking the relation between entity pairs. Consequently, relations $r_{n,(i,j)}$ and $r_{k,(i^{\prime},j^{\prime})}$ are amalgamated into a unified relation $r_{(k,n),(i,j)}=\text{LLM}(r_{n,(i,j)},{\mathbf{S}}_{n},r_{k,(i^{\prime},j^{\prime})},{\mathbf{S}}_{k},Q_{s},Q,T)$. Static segments will be associated with the entity nodes, thereby organizing the connection between the structured and long-term parts in the memory pool, as shown in Figure 2.
2.2 Reflection-based Segments Navigation
@@
Graph Expansion Search. As the number of nodes increases, the model’s ability to filter nodes and the window length both fall short of meeting the need of Entity Trial. Graph Expansion Search further leverages the structured part of the memory pool, $G$, to enhance exploration from segments. $\mathbb{E}_{s}$ is first expanded by incorporating adjacent nodes within the graph. The expansion is guided by the computation of similarity scores between the given query $Q$ and the $\mathbb{R}_{s}$, retaining nodes $\mathbb{E}^{\prime}$ with high similarity. The entity set $\mathbb{E}$ is updated through $\mathbb{E}\leftarrow\mathbb{E}^{\prime}\cup\mathbb{E}$ following each expansion iteration. The expansion process continues until either a predefined threshold of expansion iterations is reached or no new edges with sufficiently high similarity scores are found. The resultant expanded set of entities $\mathbb{E}$ is then to generate the elaborated query through ${\mathbb{Q}}=LLM(\mathbb{E},\mathbb{R},Q)$. ${\mathbb{Q}}$ is utilized to retrieve relevant segments, which are constructed into $c$ to subsequently address $Q$. $c$ is the concatenation of the original question and the generated question used to retrieve all relations.
@@
QRMeM. To trace the experience across multiple iterations, QRMeM utilizes a reflection-based navigation strategy that dynamically traverses the graph $G$, which is shown in Figure 3. Initially, QRMeM sets up an additional segmentation set, ${\mathbb{S}}_{\text{add}}$, as $\emptyset$ for further navigation. For each iteration, the real relevant segments ${\mathbb{S}}_{\text{mix}}$ is formed by combination of ${\mathbb{S}}_{\text{imp}}$ and ${\mathbb{S}}_{\text{add}}$. Then the LLM is employed to evaluate whether ${\mathbb{S}}_{\text{mix}}$ can answer the question $Q$ by the feedback of the question $Q$ and the reason $R$ for the failure is identified if not. While the ${\mathbb{S}}_{\text{mix}}$ is insufficient for the current query, the first-order adjacent nodes $\mathbb{E}_{\text{adj}}$ of the current entities and all connecting edges $\mathbb{R}$ are collected. Drawing upon the lessons learned from the previous iteration, the query, reason, and current entities $[Q;R;\mathbb{E}_{\text{s}}]$ are used to select the next expanded node. ContrieverIzacard et al. (2022) is used for calculating the similarity score between the $[Q;R;\mathbb{E}_{\text{s}}]$ and each edge in $\mathbb{R}$. After the most relevant adjust node $e$ is selected and merged with the new entity $e$ into the entity set $\mathbb{E}$, the segments corresponding to the new entity $e$ are collected. All segments not present in ${\mathbb{S}}_{\text{mix}}$ are added to ${\mathbb{S}}_{\text{add}}$ with the similarity score. Allowing for thorough updates to ${\mathbb{S}}_{\text{add}}$, a crucial step is to ensure that the context window remains sufficient after updates. ${\mathbb{S}}_{\text{add}}$ is processed and filtered according to the similarity score while adhering to the window constraints. This iterative segment combination, window checking, question evaluation, and sub-graph expansion process continues until the question $Q$ is satisfactorily answered. The detail of the algorithm is shown in Algorithm 1.
@@ 3.1 Baselines
To evaluate the effectiveness of the proposed method, we compare the QRMeM with several methods. Retrieval Methods. We use dense retrievers Contriever Izacard et al. (2022) and BM25 for all datasets. Vanilla, we use the model with 4096 context windows, including LLaMa70BChat Touvron et al. (2023) and GPT-3.5-turbo. Summary, we summarize the long context to fit the context window size. DECOMFu et al. (2021) simplifies complex questions into sub-questions, answers them sequentially. LLMLingua Pan et al. (2024), compresses prompts to speed up inference and enhance response efficiency in large language models. Long context model, including Vicuna-7b-32k, Vicuna-13b-16k, Longchat-7b-32k, Longchat-13b-16k,Llama-70b-32k. MemWalker Chen et al. (2023a), which utilize a tree of summary nodes.
@@ 3.2 Experiment Setting
The official LLaMa-70B-Chat and GPT-3.5-turbo are used for the majority of our experiments. The two models both have a maximum 4,096 token context length. We utilize the model in a zero-shot prompting fashion without further fine-tuning or using few-shot examples for our tasks. For both memory tree construction and generating action and reasoning for navigation, we use top-p sampling. The temperature is initially set to 0, but if the output cannot be parsed, we increase the temperature to 0.7 and retry up to four times. The segment size is set to 600 for all datasets.
QuALITY HotpotQA 2WikiMultihopQA MuSiQue
@@ 3.4 Main Result
1) Our study presents significant advancements over retrieval methods on the QuALITY dataset, on both the LLaMa-70B-Chat and GPT-3.5-turbo models. Detailed results related to retrieval methods are referred to in Appendix B. Notably, our approach, when compared to models utilizing longer context windows, outperforms other models designed for long-text comprehension, including GPT-3.5-16k. This highlights the effectiveness of our method in selecting relevant segments, promoting a comprehensive understanding of complex documents.
2) Furthermore, our methods demonstrated exceptional performance on Multi-Doc QA datasets. Specifically, on the 2WikiMultihopQA and MuSiQue datasets, our approach achieved remarkable results and yields comparable results to existing retrieval models on the HotpotQA dataset. Even though BM25 obtained a satisfactory result on HotpotQA, it still struggles with complex tasks like QuALITY and MuSiQue. Across all three datasets, our method outperformed long-document comprehension models. This indicates our approach’s adaptability to multi-hop tasks that require the understanding of entity relationships, capturing these relationships more effectively than long-document models. The lesser improvement on fewer inference steps HotpotQA questions and more significant advancements on the complex MuSiQue and QuALITY tasks illustrate our method’s effectiveness in organizing original paragraphs for complex tasks. GPT-3.5-16k tends to focus more on the overall context, while GPT-3.5-turbo pays closer attention to details. This is also why GPT-3.5-16k shows better performance on QuALITY datasets.
3) Three strategies proposed in our study outperformed retrieval-enhanced and direct truncation strategies. Entity Trial’s performance improvement demonstrates the effectiveness of iterative exploration. By using entities as anchors and leveraging the reflective iterative exploration of segments, it shows the impact of reflective iteration in acquiring relevant segments. Graph Expansion Search, although lacking an iterative navigation process, still utilizes structured memory to guide the LLM. This indicates that our query-based structured memory provides better guidance and relevance compared to predefined indexing. QRMeM achieves the best performance. It combines the strengths of both Entity Trial and Graph Expansion Search, using the graph to guide the model and leveraging past errors to inform future steps. This approach allows the model to utilize specific structured knowledge while retaining the ability to iteratively combine different segments.
@@ 3.7 The impact of irrelevant information
To further investigate the influence of irrelevant documents, we conducted additional tests on cross-document multi-hop questions. In this section, we retained the original wiki support text from the Long Bench dataset without chunking the long text, enabling us to more clearly identify the impact of irrelevant documents. Following Liu et al. (2023), we investigated the impact of varying the placement of supporting segments within a document, and the result is shown in Table 3. Specifically, we positioned the corresponding support segments at the beginning, middle, and tail of the overall document. Our findings are consistent with those reported by Liu et al. (2023), indicating that the placement of information can significantly affect the model’s performance on the Multi-Doc QA task. Therefore, in our subsequent experiments, we placed the multi-hop support text at both the beginning and end of the document, ensuring that the supporting segments remain in these positions to avoid any positional impact.
(a) HotpotQA
@@
Irrelevant documents utilized to increase the distance are retained, represented as [$S_{1};S_{non};S_{2}$], where $S_{1}$ and $S_{2}$ are the supporting segments, and $S_{non}$ represent sequences of non-relevant segments. The outcomes are depicted in Figure 5. The results indicate that although the model exhibits heightened attention to the positions at the beginning and end of the text, its performance is still impacted by the distance between documents. Effectiveness consistently declines as distance increases on both the HotpotQA and MuSiQue datasets.